A number is 90 and a:b=2:3 then find how much is a and b in 90.

Answer by Alan3354(69443)

A number is 90 and a:b=2:3 then find how much is a and b in 90.
--------
Pitiful wording. If you mean:
a + b = 90
and a:b=2:3
-----
a*x + b*x = 90
2x + 3x = 90
x = 18
--> a = 36, b = 54